14. Suppose a company will pay the following dividends for the next three years.

Year

Expected Dividend

1

$2

2

$3

3

$4

After the third year, the dividend will grow at a constant rate of 6% per year, the required return is 12%. What is the stock value today?
Answer #1

Year 4 dividend = 4 (1 + 6%) = 4.24

Value at year 3 = D4 / required rate - growth rate

Value at year 3 = 4.24 / 0.12 - 0.06

Value at year 3 = 4.24 / 0.06

Value at year 3 = 70.6667

Stock value = Present value of cash inflows

Stock value = 2 / (1 + 0.12)^1 + 3 / (1 + 0.12)^2 + 4 / (1 + 0.12)^3 + 70.6667 / (1 + 0.12)^3

Stock value = 1.785714 + 2.391582 + 2.847121 + 50.299161

Stock value = $57.32